# Not Every Gift Comes in Gold Paper or with a Red Ribbon: Exploring Color Perception in Text-to-Image Models

### v1

@@ D.7.1 Propriety Models Experiment
Figure S5: Propriety model performance on the prompt shown in Figure 1 (main paper). As shown, even the latest VLMs, including OpenAI 4O, Gemini 2.5, and Grok 3, often struggle to capture the correct semantics associated with multiple colors. Our method consistently outperforms existing editing approaches such as AnySD [36], FPE [18], and MasaCtrl [6], which exhibit limitations in addressing such semantic misalignment.
@@
D.7.2 Multiple Color Editing
Our editing approach can generalize to more than two objects. We create an extension of our paired-colors/objects dataset which contains three colors and objects. In this extended set, in addition to the close color we added a distant color corresponding to a third object which was chosen randomly from our object list to create prompts of the format: “a {color1} colored {object1} and a {color2} colored {object2} and a {color3} colored {object3}”. We tested our results using 3 different models which excel in multiple objects prompts: FLUX and SynGen, we added SD 2.1 too as it was our ablation model. Results are reported in Table  S7. We see that our editing approach is not limited by the number of objects, and that our method yields substantial improvements on the task of multiple color editing also over this extended set.
@@ D.7.2 Multiple Color Editing
D.7.3 Real Images Editing
We also present real image editing results using our method ColorEdit, comparing against AnySD, FPE and MasaCtrl; see Figure S7. The images were taken from the internet. As illustrated in Figure S7 we successfully edit all real images with respect to the editing prompts, including challenging objects containing fine-grained details, e.g., a bridge. We can see that ColorEdit significantly outperforms AnySD, FPE and MasCtrl in the task of real image editing.
@@ D.7.3 Real Images Editing
+FPE Refer to caption
+MasaCtrl Refer to caption
@@
+FPE Refer to caption
+MasaCtrl Refer to caption
@@
Figure S7: Above, we show edits obtained using our method ColorEdit, AnySD, FPE and MasaCtrl over real images, demonstrating that our method can modify the colors of objects in real images, outperforming baseline methods significantly.
@@ D.8 Limitations
Second, noisy cross-attention maps could also result in poor performance in some cases. While most objects are generated in distinct parts of the image, some objects are generating in overlapping regions (e.g. chair usually contains a pillow). Thus when we use our simplified prompt it struggles to correctly separate the pillow and the chair, resulting in noisy cross-attention maps that degrade the effectiveness of the Stroop attention loss. In Figure S10, we see that due to the noisy cross-attention maps the chair is colored in a similar color to the pillow. Using newer version of Stable Diffusion may potentially mitigate this gap.
@@
Figure S10: An example of a failure case due to noisy cross attention maps of semantically similar objects, e.g., a pillow on a chair which contains the pillow.

### v2

@@ D.11.1 Propriety Models Experiment
Figure S7: Propriety model performance on the prompt shown in Figure 1 (main paper). As shown, even the latest VLMs, including OpenAI 4O, Gemini 2.5, and Grok 3, often struggle to capture the correct semantics associated with multiple colors. Our method consistently outperforms existing editing approaches such as AnySD [41], MagicBrush [42], InstructPix2Pix [5], FPE [21], and MasaCtrl [7], which exhibit limitations in addressing such semantic misalignment. MagicBrush and InstructPix2Pix process the input image differently resulting scale change compare to other models with respect to grok3 image.
@@
D.11.2 Multiple Color Editing
Our editing approach can generalize to more than two objects. We create an extension of our paired-colors/objects dataset which contains three colors and objects. In this extended set, in addition to the close color we added a distant color corresponding to a third object which was chosen randomly from our object list to create prompts of the format: “a {color1} colored {object1} and a {color2} colored {object2} and a {color3} colored {object3}”. We tested our results using 3 different models which excel in multiple objects prompts: FLUX and SynGen, we added SD 2.1 too as it was our ablation model. Results are reported in Table  S12. We see that our editing approach is not limited by the number of objects, and that our method yields substantial improvements on the task of multiple color editing also over this extended set.
@@ D.11.2 Multiple Color Editing
D.11.3 Real Images Editing
We also present real image editing results using our method ColorEdit, comparing against AnySD, MagicBrush, InstructPix2Pix, FPE and MasaCtrl; see Figure S9. The images were taken from the internet. As illustrated in Figure S9 we successfully edit all real images with respect to the editing prompts, including challenging objects containing fine-grained details, e.g., a bridge. We can see that ColorEdit significantly outperforms AnySD, MagicBrush, InstructPix2Pix, FPE and MasCtrl in the task of real image editing.
@@ D.11.3 Real Images Editing
+MagicBrush Refer to caption
+InstPix2Pix Refer to caption
+FPE
+MasaCtrl
@@
+MagicBrush Refer to caption
+InstPix2Pix Refer to caption
+FPE
+MasaCtrl
@@
Figure S9: Above, we show edits obtained using our method ColorEdit, AnySD, MagicBrush, InstructPix2Pix, FPE and MasaCtrl over real images, demonstrating that our method can modify the colors of objects in real images, outperforming baseline methods significantly.
@@ D.12 Limitations
As our method optimizes the latents for performing each color edit, it is limited by the runtime required to bind the colors to the right object. We foresee future work accelerating these results, possibly using few-step diffusion models. Third, noisy cross-attention maps could also result in poor performance in some cases. Our method uses the cross attention maps extracted using the simplified prompt, while in most cases the procedure work very well, in some cases where objects are generating in overlapping regions and share mutual semantic connection (e.g. chair usually contains a pillow) our simplified prompt struggles to correctly separate the objects (e.g. the pillow from the chair), resulting noisy cross-attention maps that degrade the effectiveness of the Stroop attention loss. In Figure S12, we see that for both simplified prompt and full prompt the cross attention maps are noisy, the pillow cross-attention map is including the chair too which results edited image where the chair is colored in a similar color to the pillow.
@@
Figure S12: An example of noisy cross attention using simplified and full prompts of semantically similar objects (e.g. a pillow and a chair), this results suboptimal performance as both objects are colored with the aqua color.